OpenStudy (anonymous):

Prove: csc(2A)+cot(2A)=cot(A)

OpenStudy (anonymous):

double angles, or second powers?

OpenStudy (anonymous):

Kelly, your 2s seem ambiguous. where is the 2nd power, where is a double angle? if it is a double angle, put (2x) in parenthesis, it if is a power of 2, put a caret (^ symbol) next ti the x.

OpenStudy (anonymous):

They are double angles

OpenStudy (anonymous):

\(\large\color{black}{ \csc(2a)+\cot(2a)=\cot(a) }\)

OpenStudy (anonymous):

\(\large\color{black}{ \frac{1}{2\sin(a)\cos(a)}+\frac{\cos^2(x)-\sin^2(x)}{2\sin(a)\cos(a)}=\cot(a) }\)

OpenStudy (anonymous):

\(\large\color{black}{ \frac{1+\cos^2(x)-\sin^2(x)}{2\sin(a)\cos(a)}=\cot(a) }\)

OpenStudy (anonymous):

\(\large\color{black}{ \frac{2\cos^2(x)}{2\sin(a)\cos(a)}=\cot(a) }\)

OpenStudy (anonymous):

\(\large\color{black}{ \frac{\cos(a)}{\sin(a)}=\cot(a) }\)

OpenStudy (anonymous):

In my process whenever I said x, ignore that and regard it as a.

OpenStudy (anonymous):

apologize for my mistake, to write x and not a... I am so used to an angle of x.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

anytime... and if you have a question about any of the steps, ask.